# Efficient Classical Algorithms for Simulating Gaussian Boson Sampling on Graphs

### v1

@@ Appendix A Canonical Path Method
$\mathcal{E}_{\pi}(f,f)\geq\frac{1}{\varrho(\Gamma)}\operatorname{Var}_{\pi}f.$ (12)
To extend Theorem 8 to the multiple canonical path scenario, it suffices to prove that Lemma 9 remains valid under the generalized framework:
Proof of Lemma 9 for multiple canonical path framework.
$2\operatorname{Var}_{\pi}f$ $=\sum_{I,F\in\Omega}\pi(I)\pi(F)\left(f(I)-f(F)\right)^{2}$
$=\sum_{I,F\in\Omega}\pi(I)\pi(F)\sum_{i=1}^{l(I,F)}p^{(i)}_{IF}\cdot\left(\sum_{(M,M^{\prime})\in\gamma^{(i)}_{IF}}1\cdot\left(f(M)-f(M^{\prime})\right)\right)^{2}$
$\leq\sum_{I,F\in\Omega}\pi(I)\pi(F)\sum_{i=1}^{l(I,F)}p^{(i)}_{IF}\cdot\left|\gamma^{(i)}_{IF}\right|\sum_{(M,M^{\prime})\in\gamma^{(i)}_{IF}}\left(f(M)-f(M^{\prime})\right)^{2}$
$=\sum_{M,M^{\prime}\in\Omega}\sum_{I,F\in\Omega}\pi(I)\pi(F)\cdot\sum_{i=1}^{l(I,F)}\textbf{1}_{\{\gamma_{IF}^{(i)}\text{ uses }(M,M^{\prime})\}}p^{(i)}_{IF}\left|\gamma^{(i)}_{IF}\right|\left(f(M)-f(M^{\prime})\right)^{2}$
$=\sum_{M,M^{\prime}\in\Omega}\left(f(M)-f(M^{\prime})\right)^{2}\cdot\sum_{I,F\in\Omega}\pi(I)\pi(F)\sum_{i=1}^{l(I,F)}\textbf{1}_{\{\gamma_{IF}^{(i)}\text{ uses }(M,M^{\prime})\}}p^{(i)}_{IF}\left|\gamma^{(i)}_{IF}\right|$
$\leq\sum_{M,M^{\prime}\in\Omega}\left(f(M)-f(M^{\prime})\right)^{2}\pi(M)P(M,M^{\prime})\varrho$
$=2\varrho\mathcal{E}_{\pi}(f,f).$
Building upon the analytical framework in [16], we have extended Theorem 8 to accommodate multiple canonical paths, achieving a stronger result:
Theorem 10.
For any Markov chain, the mixing time of its lazy MC is bounded by $\tau_{\operatorname{mix}}(\varepsilon)\leq 2\varrho\left(2\ln\varepsilon^{-1}+\ln\pi_{\min}^{-1}\right)$, where $\varrho=\varrho(\Gamma)$ is the congestion defined in (10) with respect to any multiple canonical path framework $\Gamma$, and $\pi_{\min}=\min_{M\in\Omega}\pi(M)$ is the minimum value of the stationary distribution.
Appendix B Uniform Sampling for Perfect Matching
In this section, we briefly introduce algorithms for uniform sampling of perfect matchings. We consider graphs $G=(V,E)$, where $V$ is a vertex set of size $2n$ and $E$ is an edge set of size $m$. In Appendix B.1, we focus on the case of general graphs and give a MC method corresponding to Lemma 5; in Appendix B.2, we focus on the case of bipartite graphs and give a MC method corresponding to Lemma 4. Unless otherwise specified, $G$ may refer to a non-bipartite graph.
B.1 Uniform Sampling for Perfect Matchings on Dense Graphs
@@
  • If $M\in M_{n}(G)$ and $e\in M$, then move to state $M^{\prime}=M\backslash\{e\}$;
@@
  • If $M\in M_{n-1}(G)$, $u$ is matched to $w$ in $M$ and $v$ is unmatched in $M$, then move to state $M^{\prime}=M\cup\{e\}\backslash\{(w,u)\}$; symmetrically, if $v$ is matched to $w$ and $u$ is unmatched, then move to state $M^{\prime}=M\cup\{e\}\backslash\{(w,v)\}$;
@@
Note that the Markov chain is ergodic, aperiodic, and irreducible, meaning that it converges to a unique stationary distribution [17]. For two near-perfect matchings $M$ and $M^{\prime}$ such that $M^{\prime}=M\cup\{(u,v)\}\backslash\{(w,u)\}$, the corresponding transition probability is formulated as
$\Pr[M\to M^{\prime}]=\frac{1}{|E|};\quad\Pr[M^{\prime}\to M]=\frac{1}{|E|}.$
For a perfect matching $M$ and a near-perfect matching $M^{\prime}$ such that $M^{\prime}=M\backslash\{(u,v)\}$, the corresponding transition probability is formulated as
$\Pr[M\to M^{\prime}]=\frac{1}{|E|};\quad\Pr[M^{\prime}\to M]=\frac{1}{|E|}.$
@@
To bound the mixing time of this Markov chain, we introduce the concept of the conductance of a Markov chain, which is define as:
$\Phi:=\min_{S:0<\sum_{i\in S}\pi_{i}\leq\frac{1}{2}}\frac{\sum_{i\in S,j\not\in S}p_{ij}\pi_{i}}{\sum_{i\in S}\pi_{i}}$
@@
Lemma 11 ([24, 28]).
We have the following relation between the conductance $\Phi$ and the mixing time $t_{\operatorname{mix}}(\epsilon)$:
$t_{\operatorname{mix}}(\epsilon)\leq\frac{2}{\Phi^{2}}\left(\ln{\pi_{\operatorname{min}}}^{-1}+\ln\epsilon^{-1}\right).$
@@
From the above conductance bound, we obtain a polynomial bound on the mixing time of the Markov chain
Corollary 13 ([24]).
@@
$t_{\operatorname{mix}}(\epsilon)=\tilde{O}(n^{12}\ln\epsilon^{-1}).$
@@
Case 2: If $(u,v)\not\in E$, we claim that there exists $(w,z)\in M$ such that $(u,w),(v,z)\in M$. Otherwise, since $\deg(u)\geq n$ and $\deg(v)\geq n$, let $N(u):=\{w|w\text{ is the neighbor of }u\}$, $N^{\prime}(u):=\{w|(w,z)\in M\text{ and }z\text{ is the neighbor of }u\}$ and $N(v):=\{w|w\text{ is the neighbor of }v\}$. Thus we have $|N^{\prime}(u)|=|N(u)|\geq n$, $|N(v)|\geq n$ and $N^{\prime}(u)\cap N(v)=\varnothing$, but $N^{\prime}(u)\cup N(v)\subseteq V\backslash\{u,v\}$, which is a contradiction. Therefore, we can map $M$ to a perfect matching $M\cup\{(u,w),(v,z)\}\backslash\{(w,z)\}$.
For any perfect matching $M\in M_{n}(G)$, now we compute the number of near-perfect matchings $M^{\prime}\in M_{n-1}(G)$ that can be mapped to $M$ by the above scheme. For any edge $(u,v)\in M$, we can find at most one near-perfect matching $M^{\prime}\in M_{n-1}(G)$ such that $M^{\prime}=M\backslash\{(u,v)\}$ can be mapped to $M$ by case 1 of the above scheme. Thus at most $n$ near-perfect matchings $M^{\prime}\in M_{n-1}(G)$ can be mapped to $M$ by case 1 of the above scheme. Any near-perfect matching $M^{\prime}\in M_{n-1}(G)$ that can be mapped to $M$ by case 2 must be of form $M\cup\{(u,v)\}\backslash\{(u,w),(v,z)\}$. We have $\binom{n}{2}$ choices of $\{(u,w),(v,z)\}$, and each choice gives at most $4$ possible near-perfect matchings. Therefore, we can find at most $n+4\cdot\binom{n}{2}=2n^{2}-n$ near-perfect matchings $M^{\prime}\in M_{n-1}(G)$ such that $M^{\prime}$ can be mapped to $M$ by the above scheme, and we hence have
$\frac{|M_{n}(G)|}{|M_{n-1}(G)|}\geq\frac{1}{2n^{2}-n}$ $\geq\frac{1}{2n^{2}}.$
@@
By Lemma 14, the probability of getting a uniformly random perfect matching from the Markov chain is $p\geq\frac{1}{1+2n^{2}}\cdot(1-\frac{\eta}{2t})\geq\frac{1}{1+2n^{2}}\cdot\frac{1}{2}\geq\frac{1}{2+4n^{2}}$, and thus the failure probability of finally getting a uniformly random perfect matching is bounded by
@@
B.2 Uniform Sampling of Perfect Matchings on Bipartite Graphs
@@
$\Lambda(M)=\begin{cases}\lambda(M)w(u,v)&\text{ if }M\in\mathcal{M}(u,v)\text{ for some }u,v;\\ \lambda(M)&\text{ if }M\in\mathcal{M},\end{cases}$ (13)
@@
  • If $M\in\mathcal{M}$, choose an edge $e=(u,v)$ uniformly at random from $\mathcal{M}$; then move to $M^{\prime}=M\backslash\{e\}$.
@@
Thus the non-trivial transitions are of three types: removing an edge from a perfect matching; adding an edge to a near perfect matching; and exchanging an edge of a near-perfect matching with another edge adjacent to one of its holes. It is easy to verify that the stationary probabilities satisfy (13).
@@
$w^{*}(u,v)=\frac{\lambda(\mathcal{M})}{\lambda(\mathcal{M}(u,v))}$
@@
Theorem 15.
Assuming the weight function w satisfies (14) for all $(u,v)\in V_{1}\times V_{2}$ with $\mathcal{M}(u,v)\neq\varnothing$, then the mixing time of the above Markov chain is bounded above by
$O(n^{6}(\log(\pi(M)^{-1}+\log\epsilon^{-1}))).$
@@
    $\lambda_{K_{n,n}}(\mathcal{M})-1\leq$ $\lambda_{G}(\mathcal{M})\leq\lambda_{K_{n,n}}(\mathcal{M});$
    $\lambda_{K_{n,n}}(\mathcal{M}(u,v))-1\leq$ $\lambda_{G}(\mathcal{M}(u,v))\leq\lambda_{K_{n,n}}(\mathcal{M}(u,v))\text{ for }\mathcal{M}(u,v)\neq\varnothing\text{ in }G.$
@@
    $\lambda(u,v)=1\quad\text{for all pairs }(u,v).$
@@
    $w^{*}(u,v)=\frac{\lambda_{K_{n,n}}(\mathcal{M})}{\lambda_{K_{n,n}}(\mathcal{M}(u,v))}=n.$
@@
    During one iteration, we give an approximation to current $w^{*}$ within ratio $c$ for some $21<c<$ 2 for the current $\lambda$.
@@
    $\frac{w(u,v)}{w^{*}(u,v)}=\frac{\pi(\mathcal{M}(u,v))}{\pi(\mathcal{M})},$
    since $w(u,v)$ is known to us, and the probabilities on the right side may be estimated to arbitrary precision by taking several samples and take the average. Theorem 15 allows us to sample, in polynomial time, from a distribution $\hat{\pi}$ that is within variation distance $\delta$ of $\pi$.
@@
    1 Set $\hat{\eta}\leftarrow O(\eta/(n^{4}\log n))$(failure probability in each MC), $S\leftarrow O(n^{2}\log(1/\hat{\eta}))$ (number of samples in each round), $T\leftarrow O(n^{7}\log n)$ (running time of the Markov chain);
@@
Appendix C Polynomial Mixing Time for Dense Graphs
@@
Appendix C.1 proves polynomial mixing time on complete bipartite graphs. Appendix C.3 proves polynomial mixing time on complete graphs. Appendix C.2 and Appendix C.4 extend these results to dense graphs, completing the proofs of Theorem 6 and Theorem 7.
First, we provide the definitions of the canonical paths and the congestion for the matching MC. Given two matchings $I$ (initial matching) and $F$ (final matching), we need to connect $I$ and $F$ by several canonical paths $\gamma_{IF}^{(i)}$ in the adjacency graph of the matching MC, with each path $\gamma_{IF}^{(i)}$ having a corresponding probability distribution $p^{(i)}(I,F)$, which is the probability of choosing the path $\gamma_{IF}^{(i)}$ when we need to “travel” from $I$ to $F$ in the matching MC. Denote the set of canonical paths that involve the transition $t$ by $\mathrm{cp}(t):=\{(I,F,i)\mid t\in\gamma^{(i)}_{IF}\}$. Following from Theorem 10, we need to bound the congestion
ϱ:=assignitalic-$\varrho:=$ $\max_{t=(M,M^{\prime})}\bigg{\{}\frac{1}{\pi(M)P(M,M^{\prime})}\cdot\sum_{(I,F,i)\in\mathrm{cp}(t)}p^{(i)}(I,F)\cdot\pi(I)\pi(F)|\gamma_{IF}|\bigg{\}}.$ (15)
@@
C.1 Complete Bipartite Graph $G_{m,n}$
In this subsection, we provide a proof of polynomial mixing time for Algorithm 2 on the complete bipartite graph $G_{m,n}$. For convenience, suppose $m\geq n$. Unlike methods in [16], for the case of a complete graph, we can directly compute the value of the expression in (15).
Before delving into detailed analysis, we outline the proof strategy as follows. The key insight lies in leveraging the inherent symmetry of the complete graph. This symmetry allows us to establish a multiple canonical path scheme for complete bipartite graphs that symmetric transitions share identical congestion values. (It also works for complete graphs, which is shown in Appendix C.3.) This symmetry property enables us to compute the total congestion across these transitions and subsequently derive the congestion bound for one transition through direct computation.

### v2

@@ Acknowledgments
$\mathcal{E}_{\pi}(f,f)\geq\frac{1}{\varrho(\Gamma)}\operatorname{Var}_{\pi}f.$ (S6)
To extend Theorem S1 to the multiple canonical path scenario, it suffices to prove that Lemma S3 remains valid under the generalized framework:
Proof of Lemma S3 for multiple canonical path framework.
$2\operatorname{Var}_{\pi}f$ $=\sum_{I,F\in\Omega}\pi(I)\pi(F)\left(f(I)-f(F)\right)^{2}$ (S7)
$=\sum_{I,F\in\Omega}\pi(I)\pi(F)\sum_{i=1}^{l(I,F)}p^{(i)}_{IF}\cdot\left(\sum_{(M,M^{\prime})\in\gamma^{(i)}_{IF}}1\cdot\left(f(M)-f(M^{\prime})\right)\right)^{2}$ (S8)
$\leq\sum_{I,F\in\Omega}\pi(I)\pi(F)\sum_{i=1}^{l(I,F)}p^{(i)}_{IF}\cdot\left|\gamma^{(i)}_{IF}\right|\sum_{(M,M^{\prime})\in\gamma^{(i)}_{IF}}\left(f(M)-f(M^{\prime})\right)^{2}$ (S9)
$=\sum_{M,M^{\prime}\in\Omega}\sum_{I,F\in\Omega}\pi(I)\pi(F)\cdot\sum_{i=1}^{l(I,F)}\textbf{1}_{\{\gamma_{IF}^{(i)}\text{ uses }(M,M^{\prime})\}}p^{(i)}_{IF}\left|\gamma^{(i)}_{IF}\right|\left(f(M)-f(M^{\prime})\right)^{2}$ (S10)
[…]
$\leq\sum_{M,M^{\prime}\in\Omega}\left(f(M)-f(M^{\prime})\right)^{2}\pi(M)P(M,M^{\prime})\varrho$ (S12)
$=2\varrho\mathcal{E}_{\pi}(f,f).$ (S13)
Building upon the analytical framework in [18], we have extended Theorem S1 to accommodate multiple canonical paths, achieving a stronger result:
Theorem S2.
For any Markov chain, the mixing time of its lazy MC is bounded by $\tau_{\operatorname{mix}}(\varepsilon)\leq 2\varrho\left(2\ln\varepsilon^{-1}+\ln\pi_{\min}^{-1}\right)$, where $\varrho=\varrho(\Gamma)$ is the congestion defined in (S2) with respect to any multiple canonical path framework $\Gamma$, and $\pi_{\min}=\min_{M\in\Omega}\pi(M)$ is the minimum value of the stationary distribution.
Supplementary Note 2 Uniform Sampling for Perfect Matching
In this section, we briefly introduce algorithms for uniform sampling of perfect matchings. We consider graphs $G=(V,E)$, where $V$ is a vertex set of size $2n$ and $E$ is an edge set of size $m$. In Supplementary Note A, we focus on the case of general graphs and give a MC method corresponding to Lemma 2; in Supplementary Note B, we focus on the case of bipartite graphs and give a MC method corresponding to Lemma 1. Unless otherwise specified, $G$ may refer to a non-bipartite graph.
A Uniform Sampling for Perfect Matchings on Dense Graphs
@@
  • If $M\in M_{n}(G)$ and $e\in M$, then move to state $M\{$$^{\prime}=M\backslash\{e\}$;
@@
[…]
@@
[…]
$\Pr[M\to M^{\prime}]=\frac{1}{|E|};\quad\Pr[M^{\prime}\to M]=\frac{1}{|E|}.$ (S14)
[…]
$\Pr[M\to M^{\prime}]=\frac{1}{|E|};\quad\Pr[M^{\prime}\to M]=\frac{1}{|E|}.$ (S15)
@@
To bound the mixing time of this Markov chain, we introduce the concept of the conductance of a Markov chain, which is also known as bottleneck ratio:
$\Phi:=\min_{S:0<\sum_{i\in S}\pi_{i}\leq\frac{1}{2}}\frac{\sum_{i\in S,j\not\in S}p_{ij}\pi_{i}}{\sum_{i\in S}\pi_{i}}$ (S16)
@@
Lemma S4 ([31, 39]).
If the given Markov chain is irreducible, reversible, and “lazy”, i.e., $\min_{i}p_{ii}\geq\frac{1}{2}$, we have the following relation between the conductance $\Phi$ and the mixing time $t_{\operatorname{mix}}(\epsilon)$:
$t_{\operatorname{mix}}(\epsilon)\leq\frac{2}{\Phi^{2}}\left(\ln{\pi_{\operatorname{min}}}^{-1}+\ln\epsilon^{-1}\right).$ (S17)
@@
From the above conductance bound, we obtain a polynomial bound on the mixing time of the Markov chain.
Corollary S1 ([31]).
@@
$t_{\operatorname{mix}}(\epsilon)=\tilde{O}(n^{12}\ln\epsilon^{-1}).$ (S18)
@@
[…]
$\frac{|M_{n}(G)|}{|M_{n-1}(G)|}\geq\frac{1}{2n^{2}-n}$ $\geq\frac{1}{2n^{2}}.$ (S20)
@@
By Lemma S5, the probability of getting a uniformly random perfect matching from the Markov chain is $p\geq\frac{1}{1+2n^{2}}\cdot(1-\frac{\eta}{2t})\geq\frac{1}{1+2n^{2}}\cdot\frac{1}{2}\geq\frac{1}{2+4n^{2}}$, and thus the failure probability of finally getting a uniformly random perfect matching is bounded by
@@
B Uniform Sampling of Perfect Matchings on Bipartite Graphs
@@
$\Lambda(M)=\begin{cases}\lambda(M)w(u,v)&\text{ if }M\in\mathcal{M}(u,v)\text{ for some }u,v;\\ \lambda(M)&\text{ if }M\in\mathcal{M},\end{cases}$ (S21)
@@
  • If $M\in\mathcal{M}$, choose an edge $e=(u,v)$ uniformly at random from $M$; then move to $M\{$$^{\prime}=M\backslash\{e\}$.
@@
Thus the non-trivial transitions are of three types: removing an edge from a perfect matching; adding an edge to a near perfect matching; and exchanging an edge of a near-perfect matching with another edge adjacent to one of its holes. It is easy to verify that the stationary probabilities satisfy (S21).
@@
$w^{*}(u,v)=\frac{\lambda(\mathcal{M})}{\lambda(\mathcal{M}(u,v))}$ (S22)
@@
Theorem S4.
Assuming the weight function w satisfies (S23) for all $(u,v)\in V_{1}\times V_{2}$ with $\mathcal{M}(u,v)\neq\varnothing$, then the mixing time of the above Markov chain is bounded above by
$O(n^{6}(\log(\pi(M)^{-1}+\log\epsilon^{-1}))).$ (S24)
@@
    $\lambda_{K_{n,n}}(\mathcal{M})-1\leq$ $\lambda_{G}(\mathcal{M})\leq\lambda_{K_{n,n}}(\mathcal{M});$ (S25)
    $\lambda_{K_{n,n}}(\mathcal{M}(u,v))-1\leq$ $\lambda_{G}(\mathcal{M}(u,v))\leq\lambda_{K_{n,n}}(\mathcal{M}(u,v))\text{ for }\mathcal{M}(u,v)\neq\varnothing\text{ in }G.$ (S26)
@@
    $\lambda(u,v)=1\quad\text{for all pairs }(u,v).$ (S27)
@@
    $w^{*}(u,v)=\frac{\lambda_{K_{n,n}}(\mathcal{M})}{\lambda_{K_{n,n}}(\mathcal{M}(u,v))}=n.$ (S28)
@@
    During one iteration, we give an approximation to current $w^{*}$ within ratio $c$ for some $1<c<2$ for the current $\lambda$.
@@
    $\frac{w(u,v)}{w^{*}(u,v)}=\frac{\pi(\mathcal{M}(u,v))}{\pi(\mathcal{M})},$ (S29)
    since $w(u,v)$ is known to us, and the probabilities on the right side may be estimated to arbitrary precision by taking several samples and take the average. Theorem S4 allows us to sample, in polynomial time, from a distribution $\hat{\pi}$ that is within variation distance $\delta$ of $\pi$.
@@
    1 Set $\hat{\eta}\leftarrow O(\eta/(n^{4}\log n))$(failure probability in each MC), $)S\leftarrow O(n^{2}\log(1/\hat{\eta}$)) (number of samples in each round), $T\leftarrow O(n^{7}\log n)$ (running time of the Markov chain);
@@
Supplementary Note 3 Polynomial Mixing Time for Dense Graphs
@@
Supplementary Note A proves polynomial mixing time on complete bipartite graphs. Supplementary Note C proves polynomial mixing time on complete graphs. Supplementary Note B and Supplementary Note D extend these results to dense graphs, completing the proofs of Theorem 4 and Theorem 5. In Supplementary Note E we discuss the technical barriers for generating our theoretical results to non-dense graphs.
First, we provide the definitions of the canonical paths and the congestion for the matching MC. Given two matchings $I$ (initial matching) and $F$ (final matching), we need to connect $I$ and $F$ by several canonical paths $\gamma_{IF}^{(i)}$ in the adjacency graph of the matching MC, with each path $\gamma_{IF}^{(i)}$ having a corresponding probability distribution $p^{(i)}(I,F)$, which is the probability of choosing the path $\gamma_{IF}^{(i)}$ when we need to “travel” from $I$ to $F$ in the matching MC. Denote the set of canonical paths that involve the transition $t$ by $\mathrm{cp}(t):=\{(I,F,i)\mid t\in\gamma^{(i)}_{IF}\}$. Following from Theorem S2, we need to bound the congestion
$\varrho:=$ $\max_{t=(M,M^{\prime})}\bigg\{\frac{1}{\pi(M)P(M,M^{\prime})}\cdot\sum_{(I,F,i)\in\mathrm{cp}(t)}p^{(i)}(I,F)\cdot\pi(I)\pi(F)|\gamma_{IF}^{(i)}|\bigg\}.$ (S30)
@@
A Complete Bipartite Graph $G_{m,n}$
In this subsection, we provide a proof of polynomial mixing time for Algorithm 2 on the complete bipartite graph $G_{m,n}$. For convenience, suppose $m\geq n$. Unlike methods in [18], for the case of a complete graph, we can directly compute the value of the expression in (S30).
Before delving into detailed analysis, we outline the proof strategy as follows. The key insight lies in leveraging the inherent symmetry of the complete graph. This symmetry allows us to establish a multiple canonical path scheme for complete bipartite graphs that symmetric transitions share identical congestion values. (It also works for complete graphs, which is shown in Supplementary Note C.) This symmetry property enables us to compute the total congestion across these transitions and subsequently derive the congestion bound for one transition through direct computation.